A ball impinges directly on another ball at rest and is brought to rest by the impact? If halfway kinetic energy lost in the impact the coefficient of restitution is?

Grade:12th pass

1 Answers

Adarsh
6 years ago
According to question ,
½ x ½.m.u2= ½m.v2     ,where v is the velocity after collision and u is velocity before collision..
so, v2/u=2 =>v/u = 2  . ( coefficient of restitution = v/u )..